1.  Computer equipment was acquired at the beginning of year-1 at a cost of $67,000 that has an estimated residual value of $9,000 and an estimated useful life of 5 years (20 points)

  1. Determine the 1stand 2ndyear's depreciation expense, and accounts balances reported in balance sheets using the straight-line method (10 points).

Depreciation Expense:

Year 1:

Year 2:

Account Balances reported in Balance Sheet

Balance Sheets

12/31/Yr 1

12/31/Yr 2

Equipment

$

$

Accumulated Depreciation

(                                        )

(                                           )

Net Book Value of Equipment

$

$

b. Determine the 1stand 2ndyear's depreciation expense, and accounts balances reported in the balance sheets using double-declining-balance method.

Depreciation Expense:

Year 1:

Year 2:

Account Balances reported in Balance Sheet

Balance Sheets

12/31/Yr 1

12/31/Yr 2

Equipment

$

$

Accumulated Depreciation

(                                        )

(                                           )

Net Book Value of Equipment

$

$
Answer #1

Answer

  • Working

A

Cost

$            67,000.00

B

Residual Value

$              9,000.00

C=A - B

Depreciable base

$            58,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$            11,600.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

  • Requirement [a]

Depreciation Expenses:
year 1 = $ 11600
Year 2 = $ 11600

Balance Sheets

12/31/Yr 1

12/31/Yr 2

Equipment

$ 67000

$ 67000

Accumulated Depreciation

(        $ 11600                                )

(       $ 23,200                                    )

Net Book Value of Equipment

$ 55400

$ 43,800

  • Requirement [b]

Depreciation Expenses:
Year 1 = $ 67000 x 40% = $ 26,800
Year 2 = $ (67000 – 26800) x 40% = $ 16,080

Balance Sheets

12/31/Yr 1

12/31/Yr 2

Equipment

$ 67000

$ 67000

Accumulated Depreciation

(        $ 26800                             )

(       $ 42880                                    )

Net Book Value of Equipment

$ 40200

$ 24120